For every 1,000 cubic meters of air that comes through a filtering system , 0.05 gram of dust is removed.how many grams of dust are removed when 10^7 cubic meters of air have been filtered ?
(A) 5 (B) 50 (C) 500 (D) 5,000 (E) 50,000

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

For every 1,000 cubic meters of air that comes through a filtering system , 0.05 gram of dust is removed. How many grams of dust are removed when 10^7 cubic meters of air have been filtered ?
---
Proportion:
x/10^7 = 0.05/1000
---
x = 10^7(0.05/1000)
---
x = 10^4*0.05
---
Move the decimal point 4 places to the right to get:
---
x = 500
============================
Cheers,
Stan H.
============================

(A) 5 (B) 50 (C) 500 (D) 5,000 (E) 50,000

Answer by mananth(16946)

for every 10^3 cu. meters of air 0.05 gram dust
For 10^7 cu. meters. of air how much dust
10^7 *0.05 / 10^3
= 10^4 *0.05
= 500 grams of dust